General Products has an asset beta of 0.9, and a debt to equity ratio of 1.9. Their tax rate is 0.27. If the risk free rate is 0.05, and the expected return on the S&P500 is 0.09, what is the cost of levered equity for General Products?
Answer #1

Levered Beta=Asset beta*(1+(1-tax rate)*Debt/Equity)=0.9*(1+(1-0.27)*1.9)=2.15

Cost of levered equity=risk free rate+levered beta*(market return-risk free rate)=0.05+2.15*(0.09-0.05)=0.14